OpenStudy (anonymous):

(1-sinx)/cosx=cosx/(1+sinx) verify

OpenStudy (anonymous):

"cross multiply" and get \[1-\sin^2(x)=\cos^2(x)\] which is true

OpenStudy (anonymous):

but you have to prove it by changing only 1 side of the equation

OpenStudy (noelgreco):

Try multiplying numerator and denominator on the left side by the numerator's conjugate (1 + sin x)

OpenStudy (anonymous):

ok you get, 1-sinx^2/cosx(1+sinx)

OpenStudy (noelgreco):

That's right. Now look at the numerator and apply Satellite's first post.

OpenStudy (anonymous):

of cross multiplication?

OpenStudy (noelgreco):

No. We did not touch the right side, just change 1 - sin^2 x to cos^2 x and reduce.

OpenStudy (anonymous):

alright so then you get: cos^x/1+sinx, thank you sooo much!

OpenStudy (anonymous):

help on another?

OpenStudy (anonymous):

cos(pi/2+x) = -sin x I just said that they inverted the signs on both sides of the orignial identity

OpenStudy (ybarrap):

expand out cos(a + b) =cos a cos b - sin a sin b

OpenStudy (anonymous):

ok so..gimme a sec :)

OpenStudy (anonymous):

i dont think i did right...for the first product i got cos^22xpi/2...

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

hello :/

OpenStudy (anonymous):

@NoelGreco?

OpenStudy (anonymous):

unit circle?

OpenStudy (noelgreco):

The cosine of a sum is what you have to apply to the left side. ybarrap presented it correctly just substitute x for a and pi/2 for b.

OpenStudy (anonymous):

i did i got it guys

OpenStudy (ybarrap):

a=\(\pi\)/2, b = x

OpenStudy (anonymous):

thank you!

OpenStudy (ybarrap):

great!

OpenStudy (anonymous):

i frgot to refer to the unit circle to find the quantities for cos(pi/2) and sin(pi/2)

OpenStudy (ybarrap):

that helps

OpenStudy (anonymous):

yeah :P thank you both!

OpenStudy (ybarrap):

But also remember that cos and sin equal in magnitude when shifting either by a value of pi/2, so you could have figures this out by looking at the graph of sin and cos and see what happens when you shift cos(x) \(\pi\)/2 units to the left. Then you just need to know the sign polarity to make it perfect.

OpenStudy (anonymous):

ok :)

OpenStudy (noelgreco):

Also remember that if you see sin(x - pi/2) is simply cos x. That's because (x - pi/2) is simply the COmplement of x. That's where the "co" in cos, cot, csc comes from. Another example cot(x - pi/2) = tan x. Don't work harder than you have to.

OpenStudy (anonymous):

ok :)